Don is twice as old as Irene; Ed is 16 years older than Irene. The sum of all three of their ages is 60 years. Find the age of each person.
Answer by CharlesG2(834)

Don is twice as old as Irene; Ed is 16 years older than Irene. The sum of all three of their ages is 60 years. Find the age of each person.
D Don, I Irene, E Ed
D=2I
E=I+16
D+I+E=60
2I + I + I + 16 = 60
4I + 16 = 60
4I = 44
I = 11 which is Irene's age
so if Irene is 11, then Don is 22 and Ed is 27